CIRCUMSTELLAR ^{12}C/^{13}C RATIOS DERIVED FROM MILLIMETER OBSERVATIONS OF CO AND CN: INSIGHT INTO NUCLEOSYNTHESIS ON THE GIANT BRANCHES

S. N. MILAM, SETI Institute, NASA Ames Research Center, Moffett Field, CA 94035; N. J. WOOLF, \underline{L. M. ZIURYS}, University of Arizona, Steward Observatory, Department of Chemistry, Arizona Radio Observatory, Tucson, AZ 85721.

A survey of the ^{12}C/^{13}C ratio towards circumstellar envelopes has been conducted at millimeter wavelengths using the telescopes of the Arizona Radio Observatory (ARO). These ratios were obtained for a sample of local AGB, Red Giant, and Supergiant stars from observations of the ^{12}C and ^{13}C isotopologues of CO and CN. The 1 → 0 transitions of both molecules were measured at 3 mm using the 12m telescope, while the 2 → 1 lines were obtained using the Submillimeter Telescope (SMT) at 1 mm. Ratios were established from CO using radiative transfer modeling, while the hyperfine structure was used to evaluate opacity effects in CN. For C-rich shells, the ratios fall in the range ^{12}C/^{13}C ∼ 25-90, while the oxygen-rich AGB objects have values of 10-35. Ratios of ^{12}C/^{13}C ∼ 3-14 are found for the Supergiant stars. A qualitative model has been constructed based on first and third dredge-up convective mixing that can reproduce the observed ratios. Substantial mixing of H-burning (i.e. CNO) products must occur to explain the ratios in the O-rich envelopes, while a wide range of ^{12}C/^{13}C values can be generated by only a few percent mixing of He-burning ashes in the carbon-rich case.